A physics student has a single-occupancy dorm room. The student has
a small refrigerator that runs with a current of 3.00 A and a voltage of 110 V, a lamp that contains a 100-W
bulb, an overhead light with a 60-W bulb, and various other small devices adding up to 3.00 W. In Southern
California, electricity costs about 0.2 dollars per kiloWatt-hour. How much money does this student spend if
the total wattage is on for 12 hours per day for one month?
Answer #1

for the total power used

Total power used = V * I by refrigerator + Total power used by other epquipments

Total power used = 3 * 110 + 100 + 60 + 3

Total power used = 493 W

for the total energy used

total energy used = 493/1000 kW * 12 * 30 hr

total energy used = 177.5 kW.hr

total cost of energy = 177.5 kW.hr * 0.2 $/(kW.hr)

total cost of energy = 35.5 $

the total cost of energy is 35.5 $